\begin{thebibliography}{99}

\bibitem{ebe}
Belaga E. G. Evaluation of polynomials of one variable with
preliminary processing of the coefficients. Problems of
Cybernetics. {\bf 5}. Pergamon Press, 1961, 1--13.

\bibitem{ebpp}
Boyar J., Peralta R., Pochuev D. On the multiplicative complexity
of Boolean functions over the basis $(\wedge,\oplus,1)$. Theor.
Comp. Sci. 2000. {\bf 235}, 43--57.

\bibitem{eko}
Kochergin V. V. On the complexity of computations in finite
Abelian groups. Matematicheskie Voprosy Kibernetiki. 1992. {\bf
4}, 178--217.

\bibitem{elup}
Lupanov O. B. A method of circuit synthesis. Izvestiya VUZ,
Radiofizika. 1958. {\bf1}, 120--140 (in Russian).

\bibitem{en62}
Nechiporuk E. I. On the complexity of schemes in some bases
containing nontrivial elements with zero weights. Problemy
Kibernetiki. 1962. {\bf 8}, 123--160 (in Russian).

\bibitem{en63}
Nechiporuk E. I. Rectifier networks. Soviet Physics Doklady. 1963.
{\bf 8}, 5--7.

\bibitem{en69}
Nechiporuk E. I. On topological principles of self-correction.
Problemy Kibernetiki. 1969. {\bf 21}, 5--102 (in Russian).

\bibitem{epa}
Pan V. Ya. Some schemes for the evaluation of polynomials with
real coefficients. Problems of Cybernetics. {\bf 5}. Pergamon
Press, 1961, 14--32.

\bibitem{eps}
Paterson M. S., Stockmeyer L. J. On the number of nonscalar
multiplications necessary to evaluate polynomials. SIAM J. Comp.
1973. {\bf 2}, 60--66.

\bibitem{epip}
Pippenger N. The minimum number of edges in graphs with prescribed
paths. Math. System Theory. 1979. {\bf 12}, 325--346.

\bibitem{ez}
Zwick U. On the number of ANDs versus the number of ORs in
monotone Boolean circuits. Inf. Proc. Letters. 1996. {\bf 59}(1),
29--30.

\end{thebibliography}

\documentclass[12pt,a4paper]{article}
\usepackage{latexsym,amssymb,amsmath,amsthm}

\newtheorem{theorem}{Theorem}
\newtheorem{lemma}{Lemma}

\begin{document}
\title{A relation between additive and multiplicative complexity of Boolean functions\footnote{Research supported in part by RFBR, grants
11--01--00508, 11--01--00792, and OMN RAS ``Algebraic and
combinatorial methods of mathematical cybernetics and information
systems of new generation'' program (project ``Problems of optimal
synthesis of control systems'').}}
\date{}
\author{Igor S. Sergeev\footnote{e-mail: isserg@gmail.com}}

\maketitle
\begin{abstract}
In the present note we prove an asymptotically tight relation
between additive and multiplicative complexity of Boolean
functions with respect to implementation by circuits over the
basis $\{\oplus,\wedge,1\}$.
\end{abstract}

To start, consider a problem of computation of polynomials over a
semiring $(K,+,\times)$ by circuits over the arithmetic basis
$\{+,\times\} \cup K$.

It's a common knowledge that a polynomial of $n$ variables with
nonscalar multiplicative complexity $M$ (i.e. the minimal number
of multiplications to implement the polynomial, not counting
multiplications by constants) has total complexity $O(M(M+n))$.
Generally speaking, the bound could not be improved for infinite
semirings. For instance, it follows from results by E. G.
Belaga~\cite{ebe} and V. Ya. Pan~\cite{epa} (there exist
1-variable complex and real polynomials of degree $n$ with
additive complexity $n$; at the same time, each such polynomial
has nonscalar multiplicative complexity $O(\sqrt n)$~\cite{eps}).

An analogous standard bound for finite semirings is $O(M(M+n)/\log
M)$. Generally speaking, this bound is also tight in order. A
result of such sort was proven in~\cite{ez}.\footnote{\cite{ez}
deals with monotone Boolean circuits.} We prove a similar but
asymptotically tight result.

\begin{theorem}
If a Boolean function of $n$ variables can be implemented by a
circuit over the basis $\{\oplus, \wedge, 1\}$ of multiplicative
complexity $M=\Omega(n)$, then it can be implemented by a circuit
of total complexity $(1/2+o(1))M(M+2n)/\log_2 M$ over the same
basis. The bound is asymptotically optimal.
\end{theorem}

The stated result is nearly folklore, since it's an immediate
corollary of results by E. I. Nechiporuk of early 1960s. However,
these results are little known, and the corollary is even less
known. Thus, it seems appropriate to give a proof.

The second claim of the theorem (the bound optimality) holds since
almost all Boolean functions of $n$ variables have multiplicative
complexity $\sim2^{n/2}$~\cite{en62}\footnote{Instead of this
result of Nechiporuk a trivial upper bound
$\frac3{\sqrt2}\cdot2^{n/2}$ from the later paper~\cite{ebpp} is
often cited.} and total complexity $\sim2^n/n$~\cite{elup}.

Let us prove the first claim.

Let $A$ be a Boolean matrix of size $m\times n$ ($m$ rows, $n$
columns). Assign 1 to each entry of matrix which is located at
most $\log_2 m$ positions from a one of matrix $A$ in the same
row. We denote by $S(A)$ a weight\footnote{Weight of a matrix is
the number of nonzero entries in it.} of the obtained matrix and
name it an {\it active square} of matrix $A$.

The following lemma is an appropriate reformulation of particular
case of a result due to Nechiporuk~\cite{en63,en69}. In what
follows, under an implementation of a matrix we understand an
implementation of a linear operator with that matrix.

\begin{lemma}\label{1}
Any Boolean matrix $A$ of size $m\times n$ can be implemented by
an additive circuit\footnote{Over any associative and commutative
semigroup $(G, +)$.} of complexity
$\frac{S(A)}{2\log_2m}+o\left(\frac{(m+n)^2}{\log m}\right).$
\end{lemma}

\proof Divide a set of $n$ variables into groups of $s<\log_2 m$.
All possible sums in every group can be trivially computed with
complexity $<2^s$.

Regard the computed sums as new variables and note that the
problem is now reduced to implementation of a matrix of size $m
\times 2^s\lceil n/s \rceil$ and weight $\le S(A)/s$.

Divide the new matrix into horizontal sections of height $p$.
Implement each section independently. For this, in each column of
a section group all ones into pairs. Denote by $y_{i,j}$ a sum of
(new) variables corresponding to columns with paired ones from
$i$-th and $j$-th rows.

Compute all $y_{i,j}$ independently. Next, implement an $i$-th row
of a section as $y_{i,1}+\ldots+y_{i,p}+z_i$, where $z_i$ is a sum
of variables corresponding to positions with odd ones.

Note that the total complexity of computation of all $y_{i,j}$ in
all sections is at most as large as the half of matrix weight,
that is, $S(A)/(2s)$, and the number of odd ones in each section
is at most as large as the number of columns, i.e. $2^s\lceil n/s
\rceil$. Therefore, the complexity of the described circuit is
bounded from above by
$$ \frac{n2^s}{s} + \frac{S(A)}{2s} + mp + \left\lceil \frac{m}{p} \right\rceil 2^s \left\lceil \frac{n}{s} \right\rceil. $$
Assuming $p \sim m/\log^2 m$ and $s \sim \log_2 m - 3\log_2\log_2
m$, we obtain the required bound.   \qed

The bound of lemma is asymptotically tight. More general results
of that sort established by N. Pippenger~\cite{epip} and V.~V.
Kochergin~\cite{eko}.

Now we complete the proof of the theorem. Let a circuit $S$ to
implement a Boolean function $f$ with multiplicative complexity
$M$. Number all conjunction gates in the circuit in an order not
contradicting the orientation. Denote by $h_{2i-1}, h_{2i}$ input
functions of $i$-th conjunction gate, and denote by $g_i$ its
output function.

Each function $h_j$ is a linear combination of variables and
functions $g_i$, where $1\le i < j/2$. The function $f$ itself is
a linear combination of variables and all functions $g_i$.

Computation of all functions $h_j$, $j=1,\ldots,2M$, together with
the function $f$ as linear combinations of variables and functions
$g_i$ can be performed by a linear operator with matrix of size
$(2M+1)\times(M+n)$ and active square $\le (2M+1)(n+M/2+\log_2
M)$. To obtain the desired bound, implement this operator via the
method of Lemma 1.

\end{document}